# Poincaré series of semigroups

### v1

@@ Abstract.
Many invariants of affine semigroups can be studied from their Poincaré series, we offer and present several closed formulas for them. Moreover, these formulas have elementary proofs and are presented in two distinct families: one characterized by its topological nature and the other by its purely combinatorial approach.
@@ 1. Introduction
Poincaré series of one or more variables are classically used in algebra as generating functions of invariants of graded structures; see, for instance [MS], [P]. In the last 30 years, Poincaré series of several variables have been introduced in [CDK], [CDGZ1], [CDGZ2] and used to generate invariants in geometry and topology in terms of multi-index filtrations, integrations with respect Euler characteristics, and calculations with Laurent Series. Those series, say topological Poincaré series, have been computed and applied to algebraic varieties and singularities, mainly curves, surfaces and toric varieties.
For affine toric varieties, i.e. those given algebraically by toric affine semigroups, it was proved in [L] that their Poincaré series, with respect their natural filtrations given by monomial valuations, are linear images of the classical algebraic Poincaré series of affine semigroups. This poses the problem of studying the algebraic Poincaré series of the semigroup in detail and in depth and obtain formulas for it.
In this paper we study the algebraic Poincaré series of affine semigroups, namely of their semigroup algebras, following the philosophy and the techniques used in [CG], [CP] and [BCPV] for studying graded syzygies.
We offer and present several closed formulas for this Poincaré series in different terms. On one hand, topological ones, in terms of simplicial complexes associated to the semigroup elements and their versions in terms of the Alexander duals of those simplicial complexes. On the other hand, combinatorial ones, in terms of the monomial indicator series of certain key basic sets and their versions in terms of Poincaré series of colored graphs.
We also study the combinatorial impact of those formulas on complete intersections, symmetrical semigroups, and the depth of the semigroup algebra. Although in the article we have avoided syzygies, and therefore their corresponding syzygies Poincaré series in an additional variable $v$, we also show how such Poincaré series is also explicitly determined by above key sets.
@@ 2.1. Preliminaries
Along this paper $S$ will stand for an affine semigroup. This means that $S$ is a finitely generated positive cancelative commutative semigroup.
A semigroup is a set endowed with a inner operation $+$ having a 00 element. The semigroup is commutative when $+$ is commutative, it is cancelative when $m+n=m^{\prime}+n$ is only possible if $m=m^{\prime}$, it is positive when $m+n=0$ is only possible if $m=n=0$ and it is finitely generated when finitely many of his elements generate the semigroup in the sense that any element is a linear combination with non negative integers as coefficients of the generators.
@@
The cone $C(S)$ is the cone generated by the image of $S$ in $G(S)\otimes_{\mathbb{Z}}\mathbb{Q}\cong\mathbb{Q}^{d}$, i.e. the convex cone in $\mathbb{Q}^{d}$ generated by the halflines of directions $m\otimes 1$ for $m\in S$. The positive property is equivalent to $C(S)$ be a strongly convex cone, i.e. $C(S)\cap(-C(S))=\{0\}$ and $\mathbb{Q}^{d}$ is the $\mathbb{Q}$-linear space generated by $C(S)$, therefore $d$ is also the dimension of the cone $C(S)$. Moreover, $C(S)$ is a strongly convex rational polyhedral cone.
@@
Now, consider a symbol $t$ and powers $t^{m}$ for $m\in S$. The semigroup ring $R=\mathbb{Z}[S]$ is the set of expressions
@@
$a_{m}\in S$ and where there are only finitely many nonzero $a_{m}\in\mathbb{Z}$. The ring structure is given by the obvious addition and the multiplication determined by the rule $t^{m}\cdot t^{n}=t^{m+n}$.
In the same way, the power series ring $\hat{R}=\mathbb{Z}[[S]]$ is defined as above but allowing infinitely many non zero $a_{m}$ and taking into account that the rule $t^{m}\cdot t^{n}=t^{m+n}$ determines a well defined product because of the existence of the homomorphism $\lambda$, as it has finite fibers as a map. Also $\lambda$ allows to prove that the invertible series $s\in\hat{R}$ are exactly those whose coefficient $a_{0}$ (say, the independent term) is $+1$ or $-1$. In particular, $1-t^{m}$ is invertible for all $m\in S\backslash\{0\}$ and its inverse element is the series
@@
Finally, an affine semigroup is said to be a toric semigroup when the torsion subgroup $T$ of $G(S)$ is trivial, i.e. $T=0$. For the abelian group $G(S)$, viewed as a semigroup, one has larger similar objects $L=\mathbb{Z}[G(S)]$ and $\hat{L}=\mathbb{Z}[[G(S)]]$ than $R$ and $\hat{R}$ now, allowing symbols $t^{m}$ for $m\in G(S)$. Both $L$ and $\hat{L}$ are additive abelian groups. $L$ is a ring, $\hat{L}$ is not a ring but it is a $L$-module.
Toric semigroups are given, in practice, as finitely generated positive subsemigroups $S$ of $\mathbb{Z}^{h}$ for an integer $h>0$. Thus $G(S)$ is the group of lattice points of $\mathbb{Z}^{h}$ lying in the linear variety of $\mathbb{Q}^{h}$ spanned by $S$. If $d$ is the dimension of this variety, the torsion free group $G(S)$ is isomorphic to $\mathbb{Z}^{d}$, so $d$ is also the dimension of $S$.
@@
The algebraic Poincaré series of an affine semigroup $S$, or simply the Poincaré series of $S$, is the element of $\hat{R}$ given by
@@
The Poincaré series for the group $G(S)$ can be defined as the element $P_{G(S)}$ of $\hat{L}$ given by
@@ 2.2. Formulas for Poincaré series
In this subsection, we will compute the Poincaré series $P$ by two families of formulas, one in topological terms and other in combinatorial ones. Families mean that one has a formula of each type for each choice of a non empty finite subset $E$ of $S\backslash\{0\}$.
@@
Notice that $m-e_{J}$ takes sense in $G(S)$. The subsets $J$ in $T_{m}$ are called faces of dimension #$\#J-1$ of $T_{m}$. Also notice that the empty set $\emptyset$ is a face of dimension $-1$ of $T_{m}$, and that for every choice of a field $\mathbb{K}$, it is defined the $\mathbb{K}$-vector space reduced simplicial homology $\tilde{H}_{i}(T_{m})$ and cohomology $\tilde{H}^{i}(T_{m})$ both of dimension $\tilde{h}_{i}(T_{m})$, $i\geq-1$ which is 0 for $i\geq\#E-1$. The reduced homology $\tilde{h}_{i}(T_{m})$ depends om the characteristic of the field, however the reduced Euler characteristic
@@
depends only on $S$ and not on the characteristic of the field. The subcomplexes $T_{m}$ were introduced in [H] and in [CM] and later extensively in [CG], [CP], [BCPV] among others, in order to compute the syzygies of affine toric varieties. The simplicial complexes $T_{m}$ allow to use topology as their reduced homology coincides with the reduced singular homology of their topological realization in the euclidean space.
The key basic combinatorial tools we consider are the subsets $Q$, $D$ of $S$, dependent of $E$, given by
$Q$ =$=\{m\in S\,|\,m-e\notin S\quad for\,all\quad e\in E\}$
@@
$D^{J}$ =$=\{m\in S\,|\,J\subset supp(m)\quad and\quad m-n_{J}\notin S\}$
=$=\{m\in S\,|\,J\subset supp(m)\quad and\quad J\notin T_{m}\}.$
The set $D$ is the union of the sets $D^{J}$, $J\subset E$, #$\#J\geq 2$. Notice that $Q\cap D=\emptyset$, as the elements $m\in Q$ are characterized by the property $supp(m)=\emptyset$. Also they are characterized by the property $T_{m}=\{\emptyset\}$, i.e. $T_{m}$ is the simplicial complex having only the empty face.
The set $Q$ is well known in the literature and it is usually called the Apéry set with respect of $E$. The sets $D$ and $D^{J}$ were introduced in [CG].
@@
    $P=P_{S}=\dfrac{-\sum_{m\in S}\tilde{\chi}(T_{m})}{\prod_{e\in E}(1-t^{e})}$
@@
    If $q^{\prime}$, $p^{\prime}$ are other elements of $R$ such that $q^{\prime}\cdot P=p^{\prime}$ then one has $p\cdot q^{\prime}=p^{\prime}\cdot q$. The same is true in $\hat{R}$ if $p^{\prime}\in\hat{R}$.
@@
This proves (1) since the factors $1-t^{e}$ are units in $\hat{R}$.
@@
and therefore equal to the coefficient on the right.
Statement (3) follows from the algebraic fact that if $E$ generates the cone $C(S)$ then $\tilde{\chi}(T_{m})\neq 0$ for only many semigroup elements $m$. Thus the series $-\sum_{m\in S}\tilde{\chi}(T_{m})t^{m}$ is an element of $R$.
Finally (4), follows from the elementary fact
@@
the three elements $p$, $q$ and $q^{\prime}$ being in R and $p^{\prime}\in\hat{R}$. ∎
As said before, for toric semigroups the symbol $t$ is realized as a set of variables, $P$ is a power series, $p$, $q$, $p^{\prime}$ and $q^{\prime}$ are polynomials and $P=\frac{p}{q}$ is also a rational function of $\mathbb{Q}[t]$.
For general affine semigroups, q is invertible in $\hat{R}$, so $P=\frac{p}{q}=pq^{-1}$ is also a equality of elements of $\hat{R}$. If $q^{\prime}$ is invertible too, $P=\frac{p^{\prime}}{q^{\prime}}=p^{\prime}q^{\prime-1}$ as elements of $\hat{R}$ as well. For us, in practice
@@ 2.3. Particular choices of $E$
If #$\#E=1$, $E=\{e\}$, the topological tools $T_{m}$ have at most one vertex. So, there are only two possibilities: $T_{m}$ has only the empty face or it is a topological point; thus, $\tilde{\chi}(T_{m})=-1$ in the first case and $\tilde{\chi}(T_{m})=0$ in the second one. The only combinatorial tool is the Apéry set $Q=\{m\in S\leavevmode\nobreak\ :\leavevmode\nobreak\ m-e\notin S\}$, and one has $T_{m}$ is the empty face for $m\in Q$ and a point for $m\notin Q$. The Poincaré series is just given by
@@
Now, formula (1) of Poincaré series becomes shorter if one collects the terms of the sum having the same topology of $T_{m}$. Formula (2) is given by
@@
Above discussion shows how, in fact, one has many different possibilities to exhibit a fractional expression with products of cyclotomic type $1-t^{e}$. These, as corollaries of theorem 2.2 one can find nice practical consequences for the calculus.
@@
Let $E\subset S$ be a nonempty finite subset of $S$, $e_{E}=\sum_{e\in E}e$, and $Q_{E}=\{m\in S\,\,|\,\,m-e_{E}\notin S\}$ the Apery set of $S$ with respect the single element $e_{E}$. For each $J^{\prime}\subset E$, #$\#J^{\prime}\geq 1$, consider the set
@@ 3. Colored graphs
A more complete explanation of the result in section 1 can be done by means of colored graphs, a technique considered in [CG] for computing the syzygies of various graded structures of the ring $R$.
Fix a finite subset $A$ of $S$ with $A\cap E=\emptyset$ and call colors to the elements $a$ of $A$. Now, consider subsets $B$ of $S$ satisfying the property that if $b\in B$ and $n,n^{\prime}\in S$ satisfy $b+n+n^{\prime}\in B$, then $b+n\in B$ and $b+n^{\prime}\in B$. The key sets $Q$ and $D^{J}$ for $J\subset E$, #$\#E\geq 2$ satisfy that property. Then one has an oriented simple colored graph $\mathcal{G}_{B}$ defined as follows. Vertices of $\mathcal{G}_{B}$ are elements $m\in B$ such that $m-a_{I}\in B$ where $I\subset A$ and $a_{I}=\sum_{a\in I}a$. We have an oriented edge $(m,m^{\prime})$ in $\mathcal{G}_{B}$ when $m^{\prime}-m=a$ for some $a\in A$. Each edge is oriented and of color $a$. The graph is simple as $a$ is unique. The graph is infinite when $B$ is infinite. A set $I$ such that $m-a_{I}\in B$ is said to be a reason of dimension #$\#I-1$ for $m$ to be a vertex of $\mathcal{G}_{B}$.
For $-1\leq\ell\leq\#A-1$ and $m\in S$, $C_{\ell}(\mathcal{G}_{B},m)$ denote the free $\mathbb{Z}$-module generated by all the reasons of dimension $\ell$, and $\partial_{\ell}$ the composition of the simplicial boundary $\delta_{\ell}$ of the simplex $\mathcal{P}(A)$ with set of vertices $A$ with the obvious projection
@@
where $r_{\ell}(B,m)$ is the number of reasons of dimension $\ell$ for $m$ of $\mathcal{G}_{B}$, and the Poincaré series $P_{\mathcal{G}_{B}}\in\hat{R}$ of $\mathcal{G}_{B}$ given by
@@
The homology of the graphs $\mathcal{G}_{Q}$ and $\mathcal{G}_{D^{J}}$ also allows us to compare the topological data $T_{m}$ for the choice of $E$ and the corresponding simplicial complexes $T^{\prime}_{m}$ for the choice of the larger set $E^{\prime}=E\cup A$. For it, consider another colored graph $\mathcal{G}_{\bar{D}}$ on the key set $D$ defined as before but now taking as reasons of dimension $\ell$ the sets $I\cup J$, $I\subset A$, $J\subset E$, $m-a_{I}\in D^{J}$ and #$\#I\cup J=\ell+2$. Now, denote by $H_{\ell}(\mathcal{G}_{\bar{D}},m)$, $h_{\ell}(\mathcal{G}_{\bar{D}},m)$ and $\chi(\overline{D},m)$ the corresponding homology, ranks ans Euler characteristic. The Poincaré series of $\mathcal{G}_{\bar{D}}$ is defined by
@@
By adapting, without changes, the proof of theorem 2.1 in [CG] one gets
@@
Comparing with (2) in theorem 3.1 in [CG], one gets the expression of $P_{\mathcal{G}_{\bar{D}}}$ in terms of the $P_{\mathcal{G}_{D^{J}}}$

### v2

@@ Abstract.
Many invariants of finitely generated positive cancelative commutative semigroups can be studied from their Poincaré series. We offer and present several closed formulas for them. Moreover, those formulas have elementary proofs and are presented in two distinct forms: one characterized in terms of simplicial complexes and the other by a purely set theoretical approach.
Dedicated to Professors Alejandro Melle-Hernández and Enrique Artal Bartolo on the occasion of their 55th and 60th birthdays, respectively.
@@ 1. Introduction
Algebraic Poincaré series of one or more variables are classically used in algebra as generating functions of invariants of graded structures; see, for instance [MS], [P]. In the last 30 years, Poincaré series of several variables have been introduced in [CDK], [CDGZ1], [CDGZ2] and used in geometry and topology to generate invariants in terms of multi-index filtrations, integrations with respect Euler characteristics, and calculations with Laurent Series. Those series , say topological Poincaré series, have been computed and applied to algebraic varieties and singularities, mainly curves, surfaces and toric varieties.
For affine toric varieties, i.e., those given algebraically by affine semigroups, it was proved in [L] that their Poincaré series, with respect their natural filtrations given by monomial valuations, are linear images of the classical generating functions, that we also call algebraic Poincaré series of affine semigroups. This poses the problem of studying the algebraic Poincaré series of the semigroup in detail and in depth, and obtaining formulas for it.
In this paper we study the algebraic Poincaré series of finitely generated positive cancelative commutative semigroups, namely of their semigroup algebras, following the philosophy and the techniques used in [CG], [CP] and [BCPV] for studying graded syzygies.
We offer and present closed formulas for this Poincaré series in different terms. On one hand, in terms of simplicial complexes associated to the semigroup elements and their versions in terms of the Alexander duals of those simplicial complexes; on the other hand, purely set theoretical ones, in terms of the monomial indicator series of certain key basic sets and their versions in terms of colored graphs.
We also study the combinatorial impact of those formulas on complete intersections, symmetrical semigroups, and the depth of the semigroup algebra. Since they are not needed for our purposes, in the article we have avoided syzygies, and, therefore, also their corresponding syzygy Poincaré series in an additional variable $v$. However, we also discuss how such Poincaré series is also explicitly determined by above key sets.
@@ 2.1. Preliminaries
Along this paper $S$ will stand for a finitely generated positive cancelative commutative semigroup.
A semigroup is a set endowed with an inner operation $+$ having a 00 element. The semigroup is commutative when $+$ is commutative, it is cancelative when $m+n=m^{\prime}+n$ is only possible if $m=m^{\prime}$, it is positive when $m+n=0$ is only possible if $m=n=0$ and it is finitely generated when finitely many of its elements generate the semigroup in the sense that any element is a linear combination with non negative integers as coefficients of the generators.
@@
The cone $C(S)$ is the cone generated by the image of $S$ in $G(S)\otimes_{\mathbb{Z}}\mathbb{Q}\cong\mathbb{Q}^{d}$, i.e., the convex cone in $\mathbb{Q}^{d}$ generated by the halflines of directions $m\otimes 1$ for $m\in S$. The positive property is equivalent to $C(S)$ being a strongly convex cone, i.e., $C(S)\cap(-C(S))=\{0\}$ and $\mathbb{Q}^{d}$ being the $\mathbb{Q}$-linear space generated by $C(S)$, therefore $d$ is also the dimension of the cone $C(S)$. Moreover, $C(S)$ is a strongly convex rational polyhedral cone.
@@
Now, consider a symbol $t$ and another ones, called powers, $t^{m}$ for $m\in S$. The semigroup ring $R=\mathbb{Z}[S]$ is the set of expressions
@@
$a_{m}\in\mathbb{Z}$ and where there are only finitely many nonzero $a_{m}$. The ring structure is given by the obvious addition and the multiplication determined by the rule $t^{m}\cdot t^{n}=t^{m+n}$.
In the same way, the power series ring $\hat{R}=\mathbb{Z}[[S]]$ is defined as above but allowing infinitely many non zero $a_{m}$ and taking into account that the rule $t^{m}\cdot t^{n}=t^{m+n}$ determines a well defined product because of the existence of the homomorphism $\lambda$, as it has finite fibers as a map. Also $\lambda$ allows to prove that the invertible series, say the units $s$ of the ring $\hat{R}$ are exactly those whose coefficient $a_{0}$ (say, the independent term) is $+1$ or $-1$. In particular, $1-t^{m}$ is invertible for all $m\in S\backslash\{0\}$ and its inverse element is the series
@@
For the abelian group $G(S)$, viewed as a semigroup, one has larger similar objects $L=\mathbb{Z}[G(S)]$ and $\hat{L}=\mathbb{Z}[[G(S)]]$ than $R$ and $\hat{R}$ now, allowing symbols $t^{m}$ for $m\in G(S)$. Both $L$ and $\hat{L}$ are additive abelian groups. $L$ is a ring, $\hat{L}$ is not a ring but it is a $L$-module.
Finally, the semigroup is said to be an affine semigroup when the torsion subgroup $T$ of $G(S)$ is trivial, i.e., $T=0$. Affine semigroups are given, in practice, as finitely generated positive subsemigroups $S$ of $\mathbb{Z}^{h}$ for an integer $h>0$. Thus, $G(S)$ is the group of lattice points of $\mathbb{Z}^{h}$ lying in the linear variety of $\mathbb{Q}^{h}$ spanned by $S$. If $d$ is the dimension of this variety, the torsion free group $G(S)$ is isomorphic to $\mathbb{Z}^{d}$, so $d$ is also the dimension of $S$.
@@
However, general finitely generated positive cancelative commutative semigroups are not affine. A typical example for it is the semigroup $S=\ell(\mathbb{N}^{r})$, where $\ell:\mathbb{Z}^{r}\rightarrow G$ is a surjective group homomorphism, $G$ is an abelian group and $\mathbb{N}$ is the set on nonnegative integers in $\mathbb{Z}$. In this case, $G=G(S)$ is a finitely generated abelian group which has torsion in general.
Also notice that other different notions of affine semigroups appear in the literature. For instance, in [CC] affine semigroups, even topological affine semigroups are defined in the context of real topological vector spaces. But, these notions does not fit in our above context of finitely generated commutative semigroups.
@@
The algebraic Poincaré series of the semigroup $S$, or simply the generating series of $S$, is the element of $\hat{R}$ given by
@@
The algebraic Poincaré series for the group $G(S)$ can be defined as the element $P_{G(S)}$ of $\hat{L}$ given by
@@ 2.2. Formulas for Poincaré series
In this subsection, we will compute the Poincaré series $P$ by two types of formulas, one in combinatorial terms and other in set theoretical ones. It means that one has a formula of each type for each choice of a non empty finite subset $E$ of $S\backslash\{0\}$. In practice, $E$ can be any subset satisfying those properties.
@@
Notice that $m-e_{J}$ takes sense in $G(S)$. The subsets $J$ in $T_{m}$ are called faces of dimension #$\#J-1$ of $T_{m}$. Also notice that the empty set $\emptyset$ is a face of dimension $-1$ of $T_{m}$, and that for every choice of a field $\mathbb{K}$, it is defined the $\mathbb{K}$-vector space reduced simplicial homology $\tilde{H}_{i}(T_{m})$ and cohomology $\tilde{H}^{i}(T_{m})$ both of dimension $\tilde{h}_{i}(T_{m})$, $i\geq-1$ which is 0 for $i\geq\#E-1$ (see for instance [MS] page 9 for the precise definitions). The reduced homology $\tilde{h}_{i}(T_{m})$ depends on the characteristic of the field, however the reduced Euler characteristic
@@
depends only on $S$ and not on the characteristic of the field. The subcomplexes $T_{m}$ were considered and used in [H] and in [CM] and later extensively used in [CG], [CP], [BCPV] among others, in order to compute the syzygies of affine toric varieties. The simplicial complexes $T_{m}$ allow to use not only combinatorics, but also topology as their reduced homology coincides with the reduced singular homology of their topological realization in the euclidean space.
The key basic set theoretical tools we consider are the subsets $Q$, $D$ of $S$, dependent of $E$, given by
$Q$ =$=\{m\in S\,|\,m-e\notin S\text{ for all }e\in E\}$
@@
$D^{J}$ =$=\{m\in S\,|\,J\subset supp(m)\text{ and }m-e_{J}\notin S\}$
=$=\{m\in S\,|\,J\subset supp(m)\text{ and }J\notin T_{m}\}.$
The set $D$ is the union of the sets $D^{J}$, $J\subset E$, #$\#J\geq 2$. In practice, as done in [CG], this union is a good conceptual definition for $D$. Notice that $Q\cap D=\emptyset$, as the elements $m\in Q$ are characterized by the property $supp(m)=\emptyset$. Also they are characterized by the property $T_{m}=\{\emptyset\}$, i.e. $T_{m}$ is the simplicial complex having only the empty face.
The set $Q$ is well known in the literature and it is usually called the Apéry set with respect of $E$. The sets $Q$, $D$ and $D^{J}$ were introduced in [CG].
@@
    $P=P_{S}=\dfrac{-\sum_{m\in S}\tilde{\chi}(T_{m})t^{m}}{\prod_{e\in E}(1-t^{e})}$
@@
    If $q^{\prime}$, $p^{\prime}$ are elements of $R$ such that $q^{\prime}\cdot P=p^{\prime}$ then one has $p\cdot q^{\prime}=p^{\prime}\cdot q$, where $q=\prod_{e\in E}(1-t^{e})\in R=\mathbb{Z}[S]$ is a unit in $\hat{R}=\mathbb{Z}[[S]]$.
@@
This proves (1) since the factors $1-t^{e}$ are units in $\hat{R}$ and the term $\frac{1}{1-t^{e}}$ stands for the inverse series of the unit $1-t^{e}$.
@@
and therefore equal to the coefficient on the right. This coefficient is deduced to be equal by using the Newton’s binomial formula.
Statement (3) follows from the algebraic fact that if $E$ generates the cone $C(S)$ then $\tilde{\chi}(T_{m})\neq 0$ for only finitely many semigroup elements $m$. This was proved in detail in [CG]. Thus the series $-\sum_{m\in S}\tilde{\chi}(T_{m})t^{m}$ is an element of $R$.
Finally (4), follows from the elementary facts $q=\prod{e\in E}(1-t^{e})\neq 0$, it is a unit of $\hat{R}$, and
@@
the three elements $p$, $q$ and $q^{\prime}$ being in R and $p^{\prime}$ being in $\hat{R}$. ∎
As said before, for semigroups the symbol $t$ is realized as a set of variables, $P$ is a power series, $p$, $q$, $p^{\prime}$ and $q^{\prime}$ are polynomials and $P=\frac{p}{q}=\frac{p^{\prime}}{q^{\prime}}$ is also a well defined rational function of $\mathbb{Q}[t]$. This fact is elementary but, a priori, nontrivial.
For general semigroups, $q$ is invertible in $\hat{R}$, so $P=\frac{p}{q}=pq^{-1}$ is also a equality of elements of $\hat{R}$. If $q^{\prime}$ is invertible too, $P=\frac{p^{\prime}}{q^{\prime}}=p^{\prime}q^{\prime-1}$ as elements of $\hat{R}$ as well. For us, in practice
@@ 2.3. Particular choices of $E$
If #$\#E=1$, $E=\{e\}$, the combinatorial tools $T_{m}$ have at most one vertex. So, there are only two possibilities: $T_{m}$ has only the empty face or it is a topological point; thus, $\tilde{\chi}(T_{m})=-1$ in the first case and $\tilde{\chi}(T_{m})=0$ in the second one. The only set theoretical tool is the Apéry set $Q=\{m\in S\,|,\,m-e\notin S\}$, and one has $T_{m}$ is the empty face for $m\in Q$ and a point for $m\notin Q$. The algebraic Poincaré series is just given by
@@
Now, formula (1) of algebraic Poincaré series becomes shorter if one collects the terms of the sum having the same topology of $T_{m}$. Formula (2) is given by
@@
Above discussion shows how, in fact, one has many different possibilities to exhibit fractional expressions of cyclotomic polynomials of type $1-t^{e}$. Thus, as corollaries of Theorem 2.2, one can find nice practical consequences for the calculus.
@@
Let $E\subset S$ be a nonempty finite subset of $S$, $e_{E}=\sum_{e\in E}e$, and $Q_{E}=\{m\in S\,\,|\,\,m-e_{E}\notin S\}$ the Apéry set of $S$ with respect the single element $e_{E}$. For each $J^{\prime}\subset E$, #$\#J^{\prime}\geq 1$, consider the set
@@ 3. Colored graphs
A more complete explanation of the result in section 2 can be done by means of colored graphs, a technique considered in [CG] for computing the syzygies of various graded structures of the ring $R$. Here, our objective are algebraic Poincaré series, which is more modest than that of syzygies. However, we think that its explanation by means of colored graphs is interesting by itself.
Fix a finite subset $A$ of $S$ with $A\cap E=\emptyset$ and call colors to the elements $a$ of $A$. Now, consider subsets $B$ of $S$ satisfying the property that if $b\in B$ and $n,n^{\prime}\in S$ satisfy $b+n+n^{\prime}\in B$, then $b+n\in B$ and $b+n^{\prime}\in B$. The key sets $Q$ and $D^{J}$ for $J\subset E$, #$\#E\geq 2$ satisfy that property. Then one has an oriented simple colored graph $\mathcal{G}_{B}$ defined as follows. Vertices of $\mathcal{G}_{B}$ are elements $m\in B$ such that $m-a_{I}\in B$ where $I\subset A$ and $a_{I}=\sum_{a\in I}a$. We have an oriented edge $(m,m^{\prime})$ in $\mathcal{G}_{B}$ when $m^{\prime}-m=a$ for some $a\in A$. Each edge is oriented and of color $a$. The graph is simple as $a$ is unique. The graph is infinite when $B$ is infinite. A set $I$ such that $m-a_{I}\in B$ is said to be a relation (reason in [CG]) of dimension #$\#I-1$ for $m$ to be a vertex of $\mathcal{G}_{B}$.
For $-1\leq\ell\leq\#A-1$ and $m\in S$, $C_{\ell}(\mathcal{G}_{B},m)$ denote the free $\mathbb{Z}$-module generated by all the relations of dimension $\ell$, and $\partial_{\ell}$ the composition of the simplicial boundary $\delta_{\ell}$ of the full simplex $\mathcal{P}(A)$ of subsets (parts) of the set $A$, so with set of vertices $A$ with the obvious projection
@@
where $r_{\ell}(B,m)$ is the number of relations of dimension $\ell$ for $m$ of $\mathcal{G}_{B}$, and the element $P_{\mathcal{G}_{B}}\in\hat{R}$ associated to $\mathcal{G}_{B}$ given by
@@
The homology of the graphs $\mathcal{G}_{Q}$ and $\mathcal{G}_{D^{J}}$ also allows us to compare the combinatorial data $T_{m}$ for the choice of $E$ and the corresponding simplicial complexes $T^{\prime}_{m}$ for the choice of the larger set $E^{\prime}=E\cup A$. For it, consider another colored graph $\mathcal{G}_{\bar{D}}$ on the key set $D$ defined as before but now taking as relations of dimension $\ell$ the sets $I\cup J$, $I\subset A$, $J\subset E$, $m-a_{I}\in D^{J}$ and #$\#I\cup J=\ell+2$. Now, denote by $H_{\ell}(\mathcal{G}_{\bar{D}},m)$, $h_{\ell}(\mathcal{G}_{\bar{D}},m)$ and $\chi(\overline{D},m)$ the corresponding homology, ranks and Euler characteristic, the associated series to $\mathcal{G}_{\bar{D}}$ is defined by
@@
By adapting, without changes, the proof of [CG, Theorem 2.1] one gets
@@
Comparing with (2) in Theorem 3.1 with in [CG, Theorem 3.1], one gets the expression of $P_{\mathcal{G}_{\bar{D}}}$ in terms of the $P_{\mathcal{G}_{D^{J}}}$
@@
See [CG] for detailed homology computations when they are applied to homologies instead of Euler characteristics.